# Chimera: A Lossless Decoding Method for Accelerating Large Language Models Inference by Fusing all Tokens

### v1

@@
Abstract
Large language models (LLMs) have demonstrated remarkable capabilities across various tasks. However, their widespread application is hindered by the resource-intensive decoding process. To address this challenge, current approaches have incorporated additional decoding heads to enable parallel prediction of multiple subsequent tokens, thereby achieving inference acceleration. Nevertheless, the accuracy of these decoding heads falls short of the auto-regressive decoding approach.
In light of these limitations, we propose Chimera, a novel framework specifically designed for speculative sampling. Within this framework, we introduce a lightweight draft model that effectively utilizes previously generated tokens to predict subsequent words. To ensure both accuracy and efficiency, we present two strategies within the lightweight draft model. Firstly, we focus on capturing short-range dependencies at the bottom layer. Secondly, we leverage the readily available representations from the original LLM.Through empirical evaluation on the Vicuna and LlaMA-2 series, Chimera demonstrates impressive results, achieving an average latency speedup ratio of 2.7x compared to the vanilla auto-regressive decoding approach. This highlights the potential of our proposed framework in significantly improving the efficiency of large language models during the decoding process.
Chimera: A Lossless Decoding Method for Accelerating Large Language Models Inference by Fusing all Tokens
@@ 1 Introduction
Large language models (LLMs) Almazrouei et al. (2023); Touvron et al. (2023b); Achiam et al. (2023); Anil et al. (2023); Bai et al. (2023); Yang et al. (2023a) have demonstrated remarkable capabilities across diverse tasks. However, the high inference latency issue of LLMs has drawn significant attention. This latency arises partially from auto-regressive decoding, wherein the token produced at each step depends on all previously generated tokens. To alleviate the latency issue inherent to auto-regressive decoding, a novel decoding approach called speculative decoding Stern et al. (2018); Leviathan et al. (2023) has emerged.
@@
The process of speculative decoding Stern et al. (2018); Leviathan et al. (2023) can be summarized as "Draft-then-Verify." It involves two models: the original LLM and the draft model. The draft model first efficiently generates multiple speculations which consists of multiple subsequent tokens. Subsequently, the original LLM verify all of these speculations to accept the one that meets the LLM’s verification criterion. The process of drafting and verification is iterated until the termination condition is met. Given that the draft model Leviathan et al. (2023); Chen et al. (2023a) is typically smaller in size than the original LLM, it can generate multiple tokens fast, thus achieving substantial acceleration.
@@
However, the acquisition and maintenance of a separate draft model is challenging in real-world implementations. To address this issue, Blockwise Parallel Decoding Stern et al. (2018) and Medusa Cai et al. (2024) added extra decoding heads on top of the last Transformer block to predict multiple subsequent tokens in parallel. Although these methods are simple and efficient, the accuracy of the decoding heads is significantly lower than the auto-regressive decoding head, which is supported by the findings presented in Table 5. This disparity in accuracy arises from the fact that LLMs are trained to predict the next word based on all previously generated words, whereas the decoding heads predict the next word based on the prefixed sequence lacking access to the tokens generated by preceding decoding heads. Specifically, suppose the draft model is going to generate speculated sequence given the prefixed sequence , the parallel decoding head predicts the -th word based on the prefixed sequence, without having access to the tokens . The limited access to the entire context contributes to the lower accuracy observed in parallel decoding.
@@
The acceleration effect of speculative decoding primarily depends on the number of drafted tokens accepted per verification. Improving the accuracy and speed of the draft model is one of the key factors in enhancing the acceptance rate Xia et al. (2024). To enhance the accuracy of the draft model without significantly compromising speed, we propose a lightweight draft model that efficiently leverages all previously generated tokens to predict the next word.
@@
It is challenging to develop a lightweight draft model with a capacity comparable to the original LLM. We propose two strategies to cope with the challenge. The first strategy is to capture short-range dependencies instead of long-range ones at the bottom layer of the lightweight draft model. As shown in Jawahar et al. (2019); Rogers et al. (2020), lower layers of pre-trained language models mainly focus on short-range dependencies, while deeper layers are capable of capturing long-range dependencies. Only capturing short-range dependencies can reduce inference time which is highly related to the number of tokens in the input sequence. Hence, we propose a trigram encoder to encode short-range dependencies, i.e., trigrams in the input sequence. An additional advantage of the trigram encoder is that we can retrieve the hidden states of trigrams through a pre-computed lookup table, instead of computing them on the fly, further saving inference time.
@@ 2 Related Work
Independent Drafting. SpecDec Xia et al. (2023) proposed employing an independent model for the drafting process. Furthermore, Recent research has emphasized diverse knowledge distillation methods to refine small LMs into proficient drafters, enhancing behavior alignment (Miao et al., 2024; Kim et al., 2023; Zhou et al., 2023). However, this method requires a wide range of computing resources to tune the drafting model. To alleviate the computation cost, a more direct and effective method is to use the same series of small LMs as drafters to speed up inference of their larger counterparts (Leviathan et al., 2023; Spector and Re, 2023; Sun et al., 2024; Chen et al., 2023b). However, a significant behavioral gap still exists between the small LM and the target LLM, resulting in suboptimal inference accuracy.
Self-Drafting. To tackle the aforementioned challenges, certain approaches have been proposed to utilize the target LLM itself for efficient drafting (Stern et al., 2018; Santilli et al., 2023; Hooper et al., 2024). Specifically, methods like Blockwise Decoding Stern et al. (2018) and Medusa Cai et al. (2024) have introduced additional FFN heads on the top of the transformer decoder, allowing for the generation of multiple tokens simultaneously per step. Another line of research has leveraged early existing or layer-skipping techniques within the target LLM itself to handle the drafting task (Yang et al., 2023b; Zhang et al., 2023; Monea et al., 2023). For instance, Yang et al. (2023b) introduced additional subprocesses early in the current decoding step to initiate drafting future tokens in advance.
3 Methodology
In this section, we will introduce background knowledge of the auto-regressive decoding and the speculative decoding in 3.1, the lightweight draft model in 3.2, and the training details in 3.3. The general model architecture is shown in Figure 1.
@@ 3.1 Auto-regressive Decoding and Speculative Decoding
Verification. Once the drafting step is completed, all the drafted tokens are verified using the target Language Model simultaneously. The Language Model evaluates the probability or likelihood of each drafted token, enabling efficient and fast verification.
@@ 3.2.1 Trigram Encoder
We adopt a trigram encoder as the bottom layer of the lightweight draft model to save inference time without compromising too much performance. As shown in Jawahar et al. (2019); Rogers et al. (2020), lower layers of pre-trained language models mainly focus on short-range dependencies. Truncating the entire input into chunks with short lengths can reduce inference time as the computational complexity is related to the input length. For example, the time and space complexity of self-attention is quadratic in the input length Keles et al. (2023).
@@
As shown in Figure 2, the trigram encoder consists of multiple layer perceptrons (MLPs), parameterized by and respectively. Given the input is , the corresponding hidden states produced by the trigram encoder is defined as follows,
@@ 3.2.2 Full Context Encoder
We propose a full context encoder implemented via a Transformer block to capture the long-range dependencies. Capturing long-range dependencies acrosstentire context typically requires multiple Transformer blocks, but a lightweight draft model cannot afford that. Fortunately, the original LLM has properly encoded dependencies among prefix sequences. Therefore, we fully utilize the readily available hidden states from the original LLM.
@@
During training, the attention mask is meticulously designed. As long as there is a future token in a trigram, then this trigram is masked. An example illustrating the attention mask mechanism during training is presented in Figure 7 in Appendix A.
@@ 4.1 Experiment Setup
All experiments are conducted on a host system with 32 cores and 64GB of main memory, coupled with an A800 (80GB) GPU server.
Backbone LLMs. We evaluate the performance of our method on various backbone LLMs, including vicuna-7B, vicuna-13B, vicuna-33B (Chiang et al., 2023), LlaMA-2-7B, and LlaMA-2-13B (Touvron et al., 2023a). Detailed information regarding backbone LLMs can be found in Table 1.
@@
| LlaMA-2-7B | 7B | 32 | 4096 | 32 |
| LlaMA-2-13B | 13B | 40 | 5120 | 40 |
Datasets. Chimera (with specific backbone LLMs) are trained on the ShareGPT dataset (Aeala, 2022), which consists of 70,000 crafted dialogues for multi-turn conversations. The performances of all methods are evaluated in two benchmarks: MT-bench (Zheng, 2023) and Vicuna-bench (Chiang et al., 2023) to evaluate multi-turn dialogue and general capabilities, respectively.
Decoding strategies. Two decoding strategies are tested for verification: greedy decoding and typical decoding (Cai et al., 2024). For typical decoding, the temperature is set as .
@@ 4.2 Baselines
Lookahead (Fu et al., 2024). Lookahead Decoding parallelly generates multiple disjoint -grams to accommodate the future parts of a sequence. It reformulate the autoregressive decoding as the solution of nonlinear equation systems. Then, Jacobi iteration is employed for parallel decoding, capturing, validating, and integrating the generated -grams into the sequence.
Medusa-1 (Cai et al., 2024). Medusa-1 is a parameter-efficient training approach with multiple decoding heads to be trained on the same model. By relaxing the requirement for matching the distribution of the original model, Medusa enables faster and non-greedy generation and addresses the challenges of speculative decoding. Medusa-2 requires fine-tuning of the backbone LLMs, which is distinct from the approach proposed in this paper. In the subsequent sections, we will consistently refer to Medusa-1 as "Medusa" for brevity.
@@ 4.3.1 Accelearting LLM inference
Study in different decoding methods. Table 2 also test the typical decoding setting (T=1) on MT-bench, where our method exhibited an even better acceleration on greedy decoding. Specifically, the best speedup ratio is observed on Vciuna-33B, attaining a maximum speedup of 2.91x. The minimum speedup was 2.62x, with an average speedup of 2.79x, surpassing greedy search by 0.14x in terms of speed. The maximum average length was 3.38, the minimum was 3.30, and the overall average was 3.32.
@@ 4.3.2 Accurate decoding heads
the accuracy of each decoding head is in Table5 Our method outperformed Medusa with an average margin of 0.13 on Medusa head 0 and achieved a maximum improvement of 27% on top1 compared to Medusa head 0. On Medusa head 1, Medusa head 2, and Medusa head 3 , Chimera surpassed Medusa by 79%, 217%, and 267% on the top1 metric, respectively. These results demonstrate that our model exhibits higher accuracy in predicting longer sequences compared to Medusa. This also suggests that our model is better able to extract the necessary information for predicting longer sequences effectively.
@@
4.3.3 The bottleneck of inference
The time of verification vs the time of generating speculations is shown in Table 6 As the model parameters increase and the number of transformer layers grows, the Verification Time Percentage (VTP) gradually increases, reaching a maximum of 91% on Vicuna-33B, while the Generating Speculation Time Percentage (GSTP) simultaneously decreases to a minimum of 4%. This observation suggests that although our model adds a layer of transformer parameters, the investment in our model becomes more efficient in larger models, gradually reducing GSTP and yielding better returns. Moreover, the increasing dominance of VTP indicates that the time bottleneck primarily lies in the verification process. This finding also suggests that the key to further improving model speed may lie in finding faster and more effective methods for model verification.
| Model | VTP | GSTP |
| Vicuna-7B | 65% | 14% |
| Vicuna-13B | 83% | 8% |
| Vicuna-33B | 91% | 4% |
| LlaMA-2-7B | 64% | 13% |
| LlaMA-2-13B | 86% | 7% |
@@ 5 Conclusion
This paper presents Chimera, a novel framework for speculative sampling. We propose a lightweight draft model that leverages previously generated tokens to predict subsequent words. To ensure accuracy and efficiency, we introduce two strategies in the lightweight draft model: capturing short-range dependencies at the bottom layer and utilizing readily available representations from the original LLM. In experiments conducted on the Vicuna and LlaMA-2 series, Chimera achieves an average latency speedup ratio of 2.7x compared to auto-regressive decoding.
Limitations
Our method primarily focuses on improving the quality of the draft model, without optimizing the verification process. Experimental results indicate that as the model parameters increase, the proportion of verification time also gradually increases. This implies that reducing the verification time will be a crucial challenge in subsequent research on model inference acceleration.
Appendix A Attention Mask
The illustration of attention masks in full context encoder during training is shown in Figure 7.

### v2

# 5- BIG DATA MINING AND ANALYTICS 5- I S S N 22 2 0 9 6 - 0 6 54 l l 0 ? / ? ? l l p p ? ? ?– ? ? ? 5- V o l u m e 1 , N u m b e r 1, J a n u a r y 2 0 1 8

@@
Abstract: Large language models (LLMs) have demonstrated remarkable capabilities across various tasks. However, their widespread application is hindered by the resource-intensive decoding process. To address this challenge, current approaches have incorporated additional decoding heads to enable parallel prediction of multiple subsequent tokens, thereby achieving inference acceleration. Nevertheless, the accuracy of these decoding heads falls short of the auto-regressive decoding approach.In light of these limitations, we propose Chimera, a novel framework specifically designed for speculative sampling. Within this framework, we introduce a lightweight draft model that effectively utilizes previously generated tokens to predict subsequent words. To ensure both accuracy and efficiency, we present two strategies within the lightweight draft model. Firstly, we focus on capturing short-range dependencies at the bottom layer. Secondly, we leverage the readily available representations from the original LLM.Through empirical evaluation on the Vicuna and LlaMA-2-chat series, Chimera demonstrates impressive results, achieving an average latency speedup ratio of 2.7x compared to the vanilla auto-regressive decoding approach. This highlights the potential of our proposed framework in significantly improving the efficiency of large language models during the decoding process.
Key words: lossless decoding ; inference accelerating ; large language models |
6
| Ziqian Zeng, Jiahong Yu, Qianshi Pang, Huiping Zhuang, Hongen Shao, Xiaofeng Zou are with the South China University of Technology, Guangzhou, 510641, China. E-mail: zqzeng@scut.edu.cn; jiahongyugz@gmail.com; qianshipang@gmail.com; hpzhuang@scut.edu.cn; ftshaohongen@mail.scut.edu.cn; zouxiaofeng@scut.edu.cn | |
| Zihao Wang is with The Hong Kong University of Science and Technology, E-mail:zwanggc@cse.ust.hk | |
| Correspondence Author | |
| Ziqian Zeng and Jiahong Yu contribute equally to this work. | |
| Manuscript received: 2024-04-20 |
5
@@ 1 Introduction
The convergence of the cyber, physical, and social worlds has given rise to the concept of Cyber Physical Social Intelligence (CPSI), which enables the development of smart and interconnected systems across various domains. A key component of CPSI is the massive and complex Big Data that emerges from these integrated environments. Harnessing the value of this CPSI Big Data requires the employment of advanced Big Data computing techniques.
One critical challenge in CPSI Big Data computing is the need for efficient and high-performance inference of large language models (LLMs)[1, 2, 3, 4, 5, 6] , which play a pivotal role in extracting meaningful insights and powering intelligent decision-making within CPSI systems. LLMs, with their extensive knowledge and complex architectures, often face the high inference latency during inference, posing a bottleneck in the real-time or near real-time response required by CPSI applications.
@@
This latency arises partially from auto-regressive decoding, wherein the token produced at each step depends on all previously generated tokens. To alleviate the latency issue inherent to auto-regressive decoding, a novel decoding approach called speculative decoding[8, 7] has emerged.
@@
The process of speculative decoding[8, 7] can be summarized as "Draft-then-Verify." It involves two models: the original LLM and the draft model. The draft model first efficiently generates multiple speculations which consists of multiple subsequent tokens. Subsequently, the original LLM verify all of these speculations to accept the one that meets the LLM’s verification criterion. The process of drafting and verification is iterated until the termination condition is met. Given that the draft model[7, 9] is typically smaller in size than the original LLM, it can generate multiple tokens fast, thus achieving substantial acceleration.
@@
However, the acquisition and maintenance of a separate draft model is challenging in real-world implementations. To address this issue, Blockwise Parallel Decoding[8] and Medusa[10] added extra decoding heads on top of the last Transformer block to predict multiple subsequent tokens in parallel. Although these methods are simple and efficient, the accuracy of the decoding heads is significantly lower than the auto-regressive decoding head, which is supported by the findings presented in Table 5. This disparity in accuracy arises from the fact that LLMs are trained to predict the next word based on all previously generated words, whereas the decoding heads predict the next word based on the prefixed sequence lacking access to the tokens generated by preceding decoding heads. Specifically, suppose the draft model is going to generate speculated sequence given the prefixed sequence , the parallel decoding head predicts the -th word based on the prefixed sequence, without having access to the tokens . The limited access to the entire context contributes to the lower accuracy observed in parallel decoding.
@@
The acceleration effect of speculative decoding primarily depends on the number of drafted tokens accepted per verification. Improving the accuracy and speed of the draft model is one of the key factors in enhancing the acceptance rate[11]. To enhance the accuracy of the draft model without significantly compromising speed, we propose a lightweight draft model that efficiently leverages all previously generated tokens to predict the next word.
It is challenging to develop a lightweight draft model with a capacity comparable to the original LLM. We propose two strategies to cope with the challenge.
@@
The first strategy is to capture short-range dependencies instead of long-range ones at the bottom layer of the lightweight draft model. As shown in previous work[12, 13], lower layers of pre-trained language models mainly focus on short-range dependencies, while deeper layers are capable of capturing long-range dependencies. Only capturing short-range dependencies can reduce inference time which is highly related to the number of tokens in the input sequence. Hence, we propose a trigram encoder to encode short-range dependencies, i.e., trigrams in the input sequence. An additional advantage of the trigram encoder is that we can retrieve the hidden states of trigrams through a pre-computed lookup table, instead of computing them on the fly, further saving inference time.
@@ 2 Related Work
Independent Drafting. SpecDec[14] proposed employing an independent model for the drafting process. Furthermore, Recent research has emphasized diverse knowledge distillation methods to refine small LMs into proficient drafters, enhancing behavior alignment [15, 16, 17]. However, this method requires a wide range of computing resources to tune the drafting model. To alleviate the computation cost, a more direct and effective method is to use the same series of small LMs as drafters to speed up inference of their larger counterparts[7, 18, 19, 20]. However, a significant behavioral gap still exists between the small LM and the target LLM, resulting in suboptimal inference accuracy.
Self-Drafting. To tackle the aforementioned challenges, certain approaches have been proposed to utilize the target LLM itself for efficient drafting[8, 21, 22]. Specifically, methods like Blockwise Decoding[8] and Medusa[10] have introduced additional FFN heads on the top of the transformer decoder, allowing for the generation of multiple tokens simultaneously per step. Another line of research has leveraged early existing or layer-skipping techniques within the target LLM itself to handle the drafting task[23, 24, 25]. For instance[23] , introduced additional subprocesses early in the current decoding step to initiate drafting future tokens in advance.
3 Background
@@ 3.1 Auto-regressive Decoding and Speculative Decoding
Verification. Once the drafting step is completed, all the drafted tokens are verified using the target Language Model simultaneously. The target LLM evaluates the probability or likelihood of each drafted token, enabling efficient and fast verification.
@@
4 Methodology
In this section, we will introduce background knowledge of the auto-regressive decoding and the speculative decoding in 3.1, the lightweight draft model in 4.1, and the training details in 4.2. The general model architecture is shown in Figure 1.
5-
@@ 4.1.1 Trigram Encoder
We adopt a trigram encoder as the bottom layer of the lightweight draft model to save inference time without compromising too much performance. As shown in previous work[12, 13], lower layers of pre-trained language models mainly focus on short-range dependencies. Truncating the entire input into chunks with short lengths can reduce inference time as the computational complexity is related to the input length. For example, the time and space complexity of self-attention is quadratic in the input length[26].
@@
As shown in Figure 2, the trigram encoder consists of an MLP with linear layers, parameterized by and respectively. Given the input is , the corresponding hidden states produced by the trigram encoder is defined as follows,
@@ 4.1.2 Full Context Encoder
We propose a full context encoder implemented via a Transformer block to capture the long-range dependencies. Capturing long-range dependencies across the entire context typically requires multiple Transformer blocks, but a lightweight draft model cannot afford that. Fortunately, the original LLM has properly encoded dependencies among prefix sequences. Therefore, we fully utilize the readily available hidden states from the original LLM.
@@
During training, the attention mask is meticulously designed. As long as there is a future token in a trigram, then this trigram is masked. An example illustrating the attention mask mechanism during training is presented in Figure 3.
5-
@@ 4.2 Model Training
4.3 Verification
Building upon the Medusa approach[10], we employed a tree-structured attention mechanism in our verification process. Specifically, given the tree-structured draft, the original large language model computes the probability of each candidate token through a single forward pass. We then leveraged both greedy decoding[16] and typical decoding methods to sample acceptable token sequences from the draft tokens. The typical decoding approach selects candidate sequences based on their probability under the original language model, rather than using rejection sampling, thus improving efficiency without sacrificing diversity.
4.4 Inference
The inference process is described in figure 5.During the drafting step, we can sample the top-k tokens from the three chimera heads and generate a multitude of candidate sequences leveraging these tokens. We then proceed to verify these candidate sequences in parallel with the original large language model, obtaining the conditional probability of each token within the sequences. For instance, we can accept the candidate sequence with the highest overall probability through a greedy search approach.
@@ 5.1 Experiment Setup
All experiments are conducted on a server with 32-core CPU, 64 GiB host memory, and a A800(80G) GPU.
Backbone LLMs. We evaluate the performance of our method on various backbone LLMs, including Vicuna-7B, Vicuna-13B, Vicuna-33B[27], LlaMA-2-chat-7B, and LlaMA-2-chat-13B[28]. Detailed information regarding backbone LLMs can be found in Table 1.
@@
| LlaMA-2-chat-7B | 7B | 32 | 4096 | 32 |
| LlaMA-2-chat-13B | 13B | 40 | 5120 | 40 |
Datasets. Chimera (with specific backbone LLMs) are trained on the ShareGPT dataset[29], which consists of 70,000 crafted dialogues for multi-turn conversations. The performances of all methods are evaluated in two benchmarks: MT-bench[30] and Vicuna-bench[27] to evaluate multi-turn dialogue and general capabilities, respectively.
Metrics. Like other speculative sampling-based methods, Chimera primarily focuses on latency rather than throughput. We assess acceleration effects using the following metrics:
Wall-clock speedup[8]: The actual execution time acceleration ratio relative to autoregressive decoding.
Average acceptance length: The average number of tokens accepted per forward pass of the target LLM.
Top-k accuracy: The probability of i-th token predicted by the original LLM, denoted as , contained within the top-k most probable tokens predicted by the Chimera model, denoted as .
Formally, top-k accuracy of i-th token can be expressed as:
Decoding strategies. Two decoding strategies are tested for verification: greedy decoding and typical decoding[10]. For typical decoding, the temperature is set as .
@@ 5.2 Baselines
Lookahead[31]. Lookahead Decoding parallelly generates multiple disjoint -grams to accommodate the future parts of a sequence. It reformulate the autoregressive decoding as the solution of nonlinear equation systems. Then, Jacobi iteration is employed for parallel decoding, capturing, validating, and integrating the generated -grams into the sequence.
Medusa-1[10]. Medusa-1 is a parameter-efficient training approach with multiple decoding heads to be trained on the same model. By relaxing the requirement for matching the distribution of the original model, Medusa enables faster and non-greedy generation and addresses the challenges of speculative decoding.
Medusa-2 requires fine-tuning of the backbone LLMs, which is distinct from the approach proposed in this paper. In the subsequent sections, we will consistently refer to Medusa-1 as "Medusa" for brevity.
5-
@@ 5.3.1 Accelearting LLM inference
Study in different decoding methods. Table 2 also test the typical decoding setting (T=1) on MT-bench, where our method exhibited an even better acceleration on greedy decoding. Specifically, the best speedup ratio is observed on Vicuna-33B, attaining a maximum speedup of 2.91x. The minimum speedup was 2.62x, with an average speedup of 2.79x, surpassing greedy search by 0.14x in terms of speed. The maximum average length was 3.38, the minimum was 3.30, and the overall average was 3.32.
@@ 5.3.2 Accurate decoding heads
The accuracy of each decoding head is in Table 5. Our method outperformed Medusa with an average margin of 0.13 on Medusa head 0 and achieved a maximum improvement of 27% on top1 compared to Medusa head 0. On Medusa head 1, Medusa head 2, and Medusa head 3, Chimera surpassed Medusa by 79%, 217%, and 267% on the top1 metric, respectively. These results demonstrate that our model exhibits higher accuracy in predicting longer sequences compared to Medusa. This also suggests that our model is better able to extract the necessary information for predicting longer sequences effectively.
@@ 6 Conclusion
This paper presents Chimera, a novel framework for speculative sampling. We propose a lightweight draft model that leverages previously generated tokens to predict subsequent words. To ensure accuracy and efficiency, we introduce two strategies in the lightweight draft model: capturing short-range dependencies at the bottom layer and utilizing readily available representations from the original LLM. In experiments conducted on the Vicuna and LlaMA-2-chat series, Chimera achieves an average latency speedup ratio of 2.7x compared to auto-regressive decoding.
5Acknowledgement
This research was supported by the Guangzhou Basic and Applied Basic Research Foundation (Grant No. 2023A04J1687), National Natural Science Foundation of China (Grant No. 6230070401), South China University of Technology-TCL Technology Innovation Fund, the Postdoctoral Fellowship Program of CPSF (Grant No. GZC20230841).
[…]
- [1] E. Almazrouei, H. Alobeidli, A. Alshamsi, A. Cappelli, R. Cojocaru, M. Debbah, É. Goffinet, D. Hesslow, J. Launay, Q. Malartic et al., “The falcon series of open language models,” arXiv preprint arXiv:2311.16867, 2023.
- [2] H. Touvron, L. Martin, K. Stone, P. Albert, A. Almahairi, Y. Babaei, N. Bashlykov, S. Batra, P. Bhargava, S. Bhosale et al., “Llama 2: Open foundation and fine-tuned chat models,” arXiv preprint arXiv:2307.09288, 2023.
- [3] J. Achiam, S. Adler, S. Agarwal, L. Ahmad, I. Akkaya, F. L. Aleman, D. Almeida, J. Altenschmidt, S. Altman, S. Anadkat et al., “Gpt-4 technical report,” arXiv preprint arXiv:2303.08774, 2023.
- [4] R. Anil, A. M. Dai, O. Firat, M. Johnson, D. Lepikhin, A. Passos, S. Shakeri, E. Taropa, P. Bailey, Z. Chen et al., “Palm 2 technical report,” arXiv preprint arXiv:2305.10403, 2023.
- [5] J. Bai, S. Bai, Y. Chu, Z. Cui, K. Dang, X. Deng, Y. Fan, W. Ge, Y. Han, F. Huang et al., “Qwen technical report,” arXiv preprint arXiv:2309.16609, 2023.
- [6] A. Yang, B. Xiao, B. Wang, B. Zhang, C. Bian, C. Yin, C. Lv, D. Pan, D. Wang, D. Yan et al., “Baichuan 2: Open large-scale language models,” arXiv preprint arXiv:2309.10305, 2023.
- [7] Y. Leviathan, M. Kalman, and Y. Matias, “Fast inference from transformers via speculative decoding,” in ICML, vol. 202, 2023, pp. 19 274–19 286.
- [8] M. Stern, N. Shazeer, and J. Uszkoreit, “Blockwise parallel decoding for deep autoregressive models,” in NeurIPS, 2018, pp. 10 107–10 116.
- [9] C. Chen, S. Borgeaud, G. Irving, J.-B. Lespiau, L. Sifre, and J. Jumper, “Accelerating large language model decoding with speculative sampling,” 2023.
- [10] T. Cai, Y. Li, Z. Geng, H. Peng, J. D. Lee, D. Chen, and T. Dao, “Medusa: Simple llm inference acceleration framework with multiple decoding heads,” arXiv preprint arXiv: 2401.10774, 2024.
- [11] H. Xia, Z. Yang, Q. Dong, P. Wang, Y. Li, T. Ge, T. Liu, W. Li, and Z. Sui, “Unlocking efficiency in large language model inference: A comprehensive survey of speculative decoding,” arXiv preprint arXiv:2401.07851, 2024.
- [12] G. Jawahar, B. Sagot, and D. Seddah, “What does BERT learn about the structure of language?” in ACL, 2019, pp. 3651–3657.
- [13] A. Rogers, O. Kovaleva, and A. Rumshisky, “A primer in bertology: What we know about how BERT works,” TACL, vol. 8, pp. 842–866, 2020.
- [14] H. Xia, T. Ge, P. Wang, S. Chen, F. Wei, and Z. Sui, “Speculative decoding: Exploiting speculative execution for accelerating seq2seq generation,” in EMNLP, 2023, pp. 3909–3925.
- [15] X. Miao, G. Oliaro, Z. Zhang, X. Cheng, Z. Wang, Z. Zhang, R. Y. Y. Wong, A. Zhu, L. Yang, X. Shi, C. Shi, Z. Chen, D. Arfeen, R. Abhyankar, and Z. Jia, “Specinfer: Accelerating generative large language model serving with tree-based speculative inference and verification,” 2024.
- [16] S. Kim, K. Mangalam, S. Moon, J. Malik, M. W. Mahoney, A. Gholami, and K. Keutzer, “Speculative decoding with big little decoder,” in NeurIPS, 2023.
- [17] Y. Zhou, K. Lyu, A. S. Rawat, A. K. Menon, A. Rostamizadeh, S. Kumar, J.-F. Kagy, and R. Agarwal, “Distillspec: Improving speculative decoding via knowledge distillation,” 2023.
- [18] B. Spector and C. Re, “Accelerating llm inference with staged speculative decoding,” 2023.
- [19] Z. Sun, A. T. Suresh, J. H. Ro, A. Beirami, H. Jain, and F. X. Yu, “Spectr: Fast speculative decoding via optimal transport,” in NeurIPS, 2023.
- [20] Z. Chen, X. Yang, J. Lin, C. Sun, J. Huang, and K. C.-C. Chang, “Cascade speculative drafting for even faster llm inference,” 2023.
- [21] A. Santilli, S. Severino, E. Postolache, V. Maiorca, M. Mancusi, R. Marin, and E. Rodolà, “Accelerating transformer inference for translation via parallel decoding,” in ACL, 2023, pp. 12 336–12 355.
- [22] C. Hooper, S. Kim, H. Mohammadzadeh, H. Genc, K. Keutzer, A. Gholami, and S. Shao, “Speed: Speculative pipelined execution for efficient decoding,” 2024.
- [23] S. Yang, G. Lee, J. Cho, D. Papailiopoulos, and K. Lee, “Predictive pipelined decoding: A compute-latency trade-off for exact llm decoding,” 2023.
- [24] J. Zhang, J. Wang, H. Li, L. Shou, K. Chen, G. Chen, and S. Mehrotra, “Draft verify: Lossless large language model acceleration via self-speculative decoding,” 2023.
- [25] G. Monea, A. Joulin, and E. Grave, “Pass: Parallel speculative sampling,” 2023.
- [26] F. D. Keles, P. M. Wijewardena, and C. Hegde, “On the computational complexity of self-attention,” in ALT, vol. 201, 2023, pp. 597–619.
- [27] W.-L. Chiang, Z. Li, Z. Lin, Y. Sheng, Z. Wu, H. Zhang, L. Zheng, S. Zhuang, Y. Zhuang, J. E. Gonzalez, I. Stoica, and E. P. Xing, “Vicuna: An open-source chatbot impressing gpt-4 with 90%* chatgpt quality,” March 2023. [Online]. Available: https://lmsys.org/blog/2023-03-30-vicuna/
- [28] H. Touvron, L. Martin, K. Stone, P. Albert, A. Almahairi, Y. Babaei, N. Bashlykov, S. Batra, P. Bhargava, S. Bhosale, D. Bikel, L. Blecher, C. C. Ferrer, M. Chen, G. Cucurull, D. Esiobu, J. Fernandes, J. Fu, W. Fu, B. Fuller, C. Gao, V. Goswami, N. Goyal, A. Hartshorn, S. Hosseini, R. Hou, H. Inan, M. Kardas, V. Kerkez, M. Khabsa, I. Kloumann, A. Korenev, P. S. Koura, M.-A. Lachaux, T. Lavril, J. Lee, D. Liskovich, Y. Lu, Y. Mao, X. Martinet, T. Mihaylov, P. Mishra, I. Molybog, Y. Nie, A. Poulton, J. Reizenstein, R. Rungta, K. Saladi, A. Schelten, R. Silva, E. M. Smith, R. Subramanian, X. E. Tan, B. Tang, R. Taylor, A. Williams, J. X. Kuan, P. Xu, Z. Yan, I. Zarov, Y. Zhang, A. Fan, M. Kambadur, S. Narang, A. Rodriguez, R. Stojnic, S. Edunov, and T. Scialom, “Llama 2: Open foundation and fine-tuned chat models,” 2023.
- [29] Aeala, “Sharegpt_vicuna_unfiltered,” https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_unfiltered, 2022.
- [30] Q. Zheng, “Wku_nlp at semeval-2023 task 9: Translation augmented multilingual tweet intimacy analysis,” in ACL, 2023, pp. 1525–1530.
- [31] Y. Fu, P. Bailis, I. Stoica, and H. Zhang, “Break the sequential dependency of llm inference using lookahead decoding,” 2024.
- [32] X. Liu, L. Hu, P. Bailis, I. Stoica, Z. Deng, A. Cheung, and H. Zhang, “Online speculative decoding,” 2023.
5- I S S N l l 1 0 0 7 - 0 2 1 4 l l 0 ? / ? ? l l p p ? ? ?- ? ? ?
5- V o l u m e 2 2, N u m b e r 1, F e b r u a r y 2 0 1 7